The rate to rent a certain truck is $55 per day and $.20 per mile.
Your family pays $80 to rent this truck for one day. How many miles did your family drive?

Answer by ewatrrr(24785)


Hi,
Let x represent the miles driven
question states***
$55 + $0.20*x = $80
Solving for s
$0.20*x = $25
x = 125, the miles driven
CHECKING our Answer
$55 + $0.20*125 = $55 + $25 = $80